# Energy-based Contact Planning under Uncertainty for Robot Air Hockey

### v1

@@ V-B Experimental Setup
We perform 100 shots with each policy and report the accuracy score, puck speed at the goal line, and the number of bank reflections for successful shots. Each shot is initialized by placing the puck within a grid in front of the robot. Due to imperfect air flow on the air hockey table, the puck moves after release, requiring the controlled robot to adapt for a good shot.
V-C Results
Recorded metrics are reported in Table I for the simulated environment and in Table II for the real-world environment. Compared to CB and ATACOM, we observe that our framework is capable of achieving higher scoring accuracy and significantly higher puck speeds in both environments. It can be seen that different instances of our policy obtain either a high score or high puck speeds according to the corresponding parameters of the stochastic optimal control problem. For example, when compared to Ours #1, it can be seen that Ours #3 compromises scoring accuracy for faster puck speeds and a high number of bank reflections, potentially making the shots more difficult to defend against. The higher number of bank reflections produced by Ours #3 indicates that the robot kinematics allow for higher shooting speeds when hitting laterally, at the risk of missing the goal due to uncertainty gained with every bank reflection. Note that the score of Ours #3 is lower than the score chance threshold $\beta=0.5$ for this instance. This indicates that the learned model either has an error in the nominal dynamics or expects too little uncertainty gain due to bank reflections. We further note a decrease in performance for all agents due to the sim2real gap, with ATACOM showing the highest sensitivity to transfer as it requires fine-tuning on the real environment. CB shows the least decrease in performance, as it is parameterized for the real system. However, note that the shooting trajectory optimization loop of CB is slower than required to run at 50 Hz, making it prone to errors due to the puck moving unpredictably during the shooting motion. Additionally, we note higher puck speeds in the real setting for all agents as a result of differences in real and simulated contact dynamics. Examples of physical shots of all approaches can be found in the supplementary video.
VI Conclusion
This letter investigates the combination of learning-based contact planning with model-predictive robot control to achieve reactive robot behavior. We apply the proposed algorithms to the highly dynamic game of air hockey that requires long-horizon prediction and contact planning in real time. We experimentally show that our framework outperforms a purely control-based approach and a purely learning-based approach. Future work will seek to further leverage the sample efficiency of structured dynamics models to better capture the underlying contact dynamics of physical systems from real data. Furthermore, we are interested in investigating combinations of learning-based and model-based control approaches for manipulation tasks that involve more complex contact interactions.

### v2

@@ V-C Results
V-D Ablation Studies
Ablation study on reduced action space. We conduct an ablation study on the impact of learning reduced actions, i.e. learning shooting angles (1 DoF) as described in Sec. IV-B, compared to learning full actions, i.e. the mallet state at impact (4 DoF). Table III reports numerical results on the shooting performance using the same number of demonstrations for both modes. It shows that our online inference algorithm with warm-starting also works with higher-dimensional action spaces. However, a higher-dimensional action space requires more data to achieve similar performance.
Ablation study on chance constraint. We furthermore conduct a study on the impact of the threshold $\beta$, which constrains the likelihood of hitting the goal as described in Sec. IV-A. Fig. 7 illustrates how the shooting behavior changes for various $\beta$ when optimizing for puck speed ($\lambda_{1}{=}0$, $\lambda_{2}{=}1$). The score increases with $\beta$, while the main sources of uncertainty, puck speed and number of banks, are reduced with an increasing $\beta$.
VI Limitations
Several contributing factors enable the highly performative behavior displayed by our robot air hockey agent. Firstly, the dimension of the task space is low, as the puck is constrained to a plane. While this makes our results relevant for many pushing tasks, most real-world manipulation tasks involve higher-dimensional task spaces. Along the same line, we were able to leverage existing physical models as a strong prior on contact dynamics. Such models are not available for tasks that involve more complex contact dynamics. While our approach enables learning on real-world data, we exploited a physics engine to collect data for learning a compact contact model.
VII Conclusion
This paper investigated the combination of learning-based contact planning with model-predictive robot control to produce agile behavior in Robot Air Hockey. We show that distilling an optimal contact planning policy through behavior cloning effectively reduces the horizon required for lower-level trajectory optimization, enabling real-time operation. We show that the proposed approach is capable of accommodating different desired behaviors and sampling-based optimization schemes. Our results show that the proposed framework outperforms a purely control-based approach and a purely learning-based approach in simulated and real-world games of robot air hockey. Future work will seek to further leverage the sample efficiency of structured dynamics models to capture the underlying contact dynamics of physical systems from real data. Additionally, integrating physically-informed priors into the implicit model is an interesting direction for increasing the data efficiency of our approach. Lastly, we are interested in investigating the applicability of our approach to higher-dimensional task spaces with more complex contact interactions.
@@ VII Conclusion
Acknowledgments
This work was supported by the Swiss National Science Foundation (SNSF) through the CODIMAN project, by the State Secretariat for Education, Research and Innovation in Switzerland for participation in the European Commission’s Horizon Europe Program through the INTELLIMAN project (https://intelliman-project.eu/, HORIZON-CL4-Digital-Emerging Grant 101070136) and the SESTOSENSO project (http://sestosenso.eu/, HORIZON-CL4-Digital-Emerging Grant 101070310), by the China Scholarship Council (Grant 201908080039) and by the German Federal Ministry of Education and Research (BMBF) through the KIARA project (Grant 13N16274).
TABLE III: Reduced Action Space (R. Act.) v. Full Action Space (F. Act.).
R. Act. 0.80 1.44 $\pm$ 0.63 0.53
F. Act. 0.51 0.96 $\pm$ 0.13 0.00
Figure 7: Impact of the chance constraint threshold $\beta$ on shooting performance.